# Survey of Computerized Adaptive Testing: A Machine Learning Perspective

### 05 Apr 2024 v2

@@ 9. Opportunities for Future Research
Future research should harness machine learning to enhance the multi-dimensionality of the assessment. This involves not only the traditional response patterns but also the nuanced analysis of process data such as response times and mouse movements, which can provide insights into an examinee’s problem-solving strategies and levels of engagement. Moreover, integrating learning data, such as the examinee’s prior interactions with materials, can offer a longitudinal perspective on their learning trajectory and readiness for new concepts. Additionally, the analysis of educational content, encompassing textual, visual, and auditory materials (Xiao et al., 2022; Shi et al., 2014), allows for a richer understanding of how examinees interact with multifaceted information. Such machine learning-driven approaches promise to refine CAT systems comprehensively, enabling them to deliver assessments that are not just accurate reflections of an examinee’s proficiency but also predictive of their potential for future learning.
Multi-Stage Testing
Multi-Stage Testing (MST) is a variant of CAT that is frequently employed in high-stakes examinations (Magis et al., 2017). Unlike CAT, which select a single question at each step, MST presents a set of questions (i.e., testlet) at each step, which is pre-designed by experts as the unit of selection. The selection process is predetermined by the test design, based on a decision tree or a similar algorithmic structure, aiming to match test takers’ estimated proficiency levels more closely (Mead, 2006). Building upon CAT, MST presents more complex challenges, including considerations around the number of testlets, the size of each testlet, and questions’ combinatory effects within a testlet. Currently, the selection algorithm for MST relies heavily on manual design. To facilitate the widespread adoption of MST for large-scale testing, there is a need for increased research into the automated MST construction.
Empowering CAT with Generative AI
Generative artificial intelligence, with Large Language Models (LLMs) at the forefront, is trained on massive, cross-domain datasets, endowing it with versatility and a profound repository of world knowledge (Dong et al., 2023; Chang et al., 2023). These models have already shown preliminary progress in user modeling, such as recommendation systems, and in the generation of personalized strategy. In the future, there is potential for these large models to significantly enhance CAT systems in various aspects, such as question selection, proficiency assessment, and even the automatic generation of novel, tailored questions on the fly (Bhandari et al., 2023) – questions that are not pre-existing in the bank. This could lead to more effective and individualized testing experiences.
Towards Explainable Machine Learning in CAT
@@
Traditional CAT systems, particularly those based on information and statistic approaches, are lauded for their interpretability, from the parameters of CDMs to the logic behind the question selection algorithm. This transparency provides valuable insights to all stakeholders, including examinees, parents, and educators, and supports developers in debugging and refining the CAT system behind the final testing scores. However, recent machine learning approaches, especially those employing deep learning, have an overwhelming advantage in capabilities on knowledge discovery, at the cost of reduced interpretability. Bridging the gap between these paradigms to create CAT systems that are both accurate and self-explanatory is a significant challenge that future research must address. This is particularly crucial for high-stakes standardized testing, where the outcomes carry significant consequences
CAT for AI System Evaluation
Computerized Adaptive Testing (CAT) has traditionally been tailored for educational purposes, primarily serving humans. However, the intrinsic qualities of CAT—its efficiency at discerning test-taker’s cognitive states and its capacity for conducting multi-step, adaptive interactions—suggest a broader potential. Given these capabilities, we can foresee its adaptation to not only serve a wider human audience but also to assess and refine AI systems. In the domain of AI, CAT methodologies could become crucial in the efficient evaluation of machine learning models, including their cognitive capabilities. This is particularly relevant in the advanced Large Language Models (LLMs) that are approaching human-level performance, where CAT could provide nuanced insights into their cognitive-like functions. Moreover, CAT could expedite model capacity evaluation using the least possible amount of samples. This evolution can deepen our understanding of both human and machine cognition, paving the way for innovative uses in efficient evaluation services, and smarter, more responsive AI systems.
10. Conclusion
@@
This comprehensive survey underscored the intricate and vast system that is Computerized Adaptive Testing (CAT), emphasizing the potential and prospects of applying machine learning to enhance CAT. The paper focused on CAT’s two primary concerns of accuracy and efficiency within the assessment process from the machine learning perspective. The presented insights were accessible and relevant not just for specialists in education or psychometrics but for a wide array of researchers. We encourage interested readers to learn about the transformative impact of machine learning in this field and take this survey as a reference for future research.

### 09 Mar 2026 v3

@@ 9 Opportunities for Future Research
Multi-Dimensionality of the Assessment Process. Future research should harness machine learning to enhance the multi-dimensionality of the assessment. This involves not only the traditional response patterns but also the nuanced analysis of process data such as response times and mouse movements, which can provide insights into an examinee’s problem-solving strategies and levels of engagement. Moreover, integrating learning data, such as the examinee’s prior interactions with materials, can offer a longitudinal perspective on their learning trajectory and readiness for new concepts. Additionally, the analysis of content, encompassing textual, visual, and auditory materials [174, 143], allows for a richer understanding of how examinees interact with multifaceted information. Such machine learning-driven approaches promise to refine CAT systems comprehensively, enabling them to deliver assessments that are not just accurate reflections of an examinee’s proficiency but also predictive of their potential for future learning.
@@
Towards Explainable Machine Learning in CAT. Traditional CAT systems, particularly those based on information and statistic approaches, are lauded for their interpretability, from the parameters of measurement models to the logic behind the question selection algorithm. This transparency provides valuable insights to all stakeholders, including examinees, parents, and educators, and supports developers in debugging and refining the CAT system. In practice, interpretability is often important for deploying CAT in high-stakes settings: test providers may need to explain why certain items were chosen, demonstrate fair treatment across groups, and support audits or appeals. Even if the selection policy is complex, this can be addressed by making the main constraints transparent and recording a simple, human-readable reason for each selection. However, recent machine learning approaches, especially those employing deep learning, have an overwhelming advantage in capabilities on knowledge discovery, at the cost of reduced interpretability. Bridging the gap between these paradigms to create CAT systems that are both accurate and self-explanatory is a significant challenge that future research must address. This is particularly crucial for high-stakes standardized testing, where the outcomes carry significant consequences.
Empowering CAT with Generative AI. Generative artificial intelligence (e.g., LLMs) is trained on massive, cross-domain datasets, endowing it with versatility and a profound repository of world knowledge [44, 27]. These models have already shown preliminary progress in user modeling, such as recommendation systems, and in the generation of personalized strategy [176, 187]. This connection is conceptually aligned with CAT: both aim to infer latent user traits (e.g., proficiency) from observed behavior and then adapt subsequent interactions accordingly. LLMs/agents can enrich the observation space beyond binary correctness by leveraging intermediate steps, explanations, hesitation patterns, and error types, which can support finer-grained proficiency estimation when properly calibrated. In the future, there is potential for these large models to significantly enhance CAT systems in various aspects, such as question selection, proficiency assessment, and even the automatic generation of novel, tailored questions on the fly [15, 132] – questions that are not pre-existing in the bank. A practical integration is to use LLMs as assistive modules: they can draft candidate items conditioned on a targeted construct/skill label, format constraints, and an intended difficulty region, and produce useful metadata (topic tags, expected solution outlines, and common misconceptions) that helps index and retrieve items efficiently. We can envision a future where testing paradigms evolve towards greater intelligence and automation. A well-trained testing agent could engage with examinees in natural language interactions, utilizing various cues and process details to conduct a comprehensive assessment of abilities. This approach would move beyond the monotonous task of having examinees respond to questions from a predefined bank or benchmark one by one. Such advancements could lead to more effective and personalized testing experiences.
Improving Machine Intelligence Evaluation. Traditional AI model evaluation relies extensively on large, gold-standard benchmarks. The maxim “more is better” has driven the use of larger benchmarks to provide comprehensive assessments. However, the sheer size of these benchmarks incurs significant time and computational costs, making fast and economical evaluations challenging. For example, evaluating the performance of a single LLM on the full HELM benchmark can consume over 4,000 GPU hours (or cost over \$10,000 for APIs) [99]. Moreover, these benchmarks are often plagued by low-quality questions, errors, and contamination issues [133]. As discussed, an increasing number of researchers are attempting to leverage CAT and psychometrics to identify and address these issues, reducing evaluation overhead and gradually transforming it into a new evaluation paradigm. This shift is especially valuable as AI systems approach human-level performance, where CAT can offer finer-grained analysis of cognitive-like behaviors. Although advanced LLMs differ fundamentally from humans in architecture, their learned behaviors often exhibit similar characteristics, since they are trained on large-scale human-produced data and display cognitive-like signatures[190]. CAT does not assume models are “human”; it only requires observable responses that can be consistently scored and related to item statistics. Ultimately, this emerging paradigm may lead to smarter, faster, and more cost-effective evaluations: deepening our understanding of both human and machine intelligence.
10 Conclusion
[…]
@@ 10 Conclusion
This comprehensive survey has highlighted the intricate and expansive nature of CAT, emphasizing the potential and prospects of integrating machine learning to enhance CAT systems. The paper primarily focused on the dual concerns of accuracy and efficiency within machine/human assessment. The insights presented are accessible and relevant not only to specialists in education and psychometrics but also to a broad spectrum of researchers. We encourage interested readers to explore the transformative impact of machine learning in this field and to use this survey as a reference for future research.